# Bayesian spatiotemporal modelling of political violence and conflict events using discrete-time Hawkes processes

### v1

@@ Abstract
The monitoring of conflict risk in the humanitarian sector is largely based on simple historic averages. To advance our understanding, we propose Hawkes processes, a self-exciting stochastic process used to describe phenomena whereby past events increase the probability of future events occurring. The overarching goal of this work is to assess the potential for using a more statistically rigorous approach to monitor the risk of political violence and conflict events in practice, and characterise their temporal and spatial patterns.
The region of South Asia was selected as an exemplar of how our model can be applied globally. We individually analyse the various types of conflict events for the countries in this region. In particular, a Bayesian, spatiotemporal variant of the Hawkes process is fitted to data gathered by the Armed Conflict Location and Event Data (ACLED) project to obtain sub-national estimates of conflict risk over time and space. We find that our model can effectively estimate the risk level of these events within a statistically sound framework, with a more precise understanding of the uncertainty around these estimates than was previously possible. The model also provides insights into differences in behaviours between countries and conflict types. This enables a better understanding of the temporal and spatial patterns inherent in the data which can help to inform preventative measures.
We demonstrate the advantages of the Bayesian framework by comparing our results to a less computationally intensive approach using maximum likelihood estimation. We find that while the maximum likelihood approach gives reasonable point estimates, the Bayesian approach is preferred when possible. Practical examples are presented to demonstrate how the proposed model can be used to monitor conflict risk. Comparing to current practices that rely on historical averages, we also show that our model is more stable and robust to outliers. In showcasing these promising results, we aim to support actors in the humanitarian and political sciences sectors in making data-informed decisions, such as the allocation of resources in conflict-prone regions.
@@ 1 Introduction
Hawkes processes (Hawkes, 1971) are self-exciting stochastic processes whereby previous events in the process increase the probability of future events occurring. They have previously been used to study phenomena such as earthquakes (Ogata, 1988), crime and terrorism (White et al., 2013; Mohler et al., 2011), and social media interactions (Chen and Tan, 2018; Dutta et al., 2020). In this article, we extend the standard temporal model to include a spatial kernel that quantifies the spatial impact of events to the intensity of the process. A discrete-time variant of the spatiotemporal Hawkes process is fit to the ACLED data at a fine spatial scale to obtain sub-national estimates of conflict risk. This approach aids in characterising the temporal and spatial patterns that emerge from these types of conflicts, and provides an estimate of risk over space and time that can be used both for real-time monitoring and for making short-term predictions. We use data from the region of South Asia to showcase our approach; however, the methodology can easily be extended to other countries. For the purposes of this example, the four countries considered are Bangladesh, Sri Lanka, Nepal and Pakistan. In addition to modelling each of the countries separately, the various types of conflict events in each country are also modelled individually. By doing so, we can compare the spatial and temporal dynamics of events between both countries and conflict types.
@@
1.1 Discrete-time Hawkes processes
Hawkes processes (HPs) (Hawkes, 1971) were introduced to model phenomena that exhibit self-exciting behaviour whereby events occur in groups or clusters. HPs were originally introduced as a continuous-time process. However, since the data collected by the ACLED project is in the form of daily counts we adopt a discrete-time version of the HP (Linderman and Adams, 2015; Browning et al., 2021; Mohler, 2013), referred to in this article as the discrete-time Hawkes process (DTHP).
For a given DTHP we assume that the number of events that occur at a particular time is distributed according to a Poisson distribution. This distribution can be characterised by its conditional intensity function which governs the rate at which events occur. The intensity function is comprised of two components: the baseline process and the self-exciting process. The baseline process describes the arrival rate of events that occur independently, and not as a result of other events. The self-exciting component on the other hand, describes the clustering behaviours of the process. The aspect of the self-exciting component that describes the impact that past events have on the current intensity of the process is referred to as the triggering kernel, and is a function of the time that has elapsed since an event occurred.
To state this formally, consider a linear DTHP $N$, where $N(t)$ represents the number of events up to the $t$-th time interval. $N(t)$ is dependent on the set of events to occur up to but not including time $t$, denoted by $\mathcal{H}_{t-1}=\{y_{s}:s\leq t-1\}$, where $y_{s}$ represents the observed number of events during a given time interval $s$. We denote the conditional intensity function as $\lambda(t)$. This quantity represents the number of events expected to occur in the counting process $N(t)$ during time interval $t$, conditionally on the past. Furthermore, the random variable $Y(t)=N(t)-N(t-1)$ represents the number of event occurrences in time interval $(t-1,t]$, and thus,
$\lambda(t)$ $=\mathbb{E}\{N(t)-N(t-1)|\mathcal{H}_{t-1}\}$
$=\mu+\alpha\sum_{i:t_{i}<t}y_{t_{i}}g(t-t_{i})$ (1)
where $\mu$ represents the baseline mean of the process and the second term represents the self-exciting component of the HP. In particular, $\alpha$ is a magnitude parameter that describes the expected number of events that a single event is expected to produce. The triggering kernel $g(t-t_{i})$ is a probability mass function that describes the influence of these past events on the intensity of the process, given the time elapsed since time $t_{i}$, where $t>t_{i}$.
The number of events observed during time interval $t$, namely $y_{t}$, is distributed according to the random variable, $Y(t)$, which has conditional mean $\mathbb{E}\{Y(t|\mathcal{H}_{t-1})\}=\lambda(t)$ as defined in (1). In this analysis, $Y(t)$ is assumed to be Poisson distributed such that $Y(t|\mathcal{H}_{t-1})\sim\text{Pois}(\lambda(t))$. Thus, for the proposed DTHP model, the probability that time interval $t$ has $y_{t}$ events is,
$P(Y(t)=y_{t}|\mu,\alpha,g)=\frac{(\mu+\alpha\sum_{i:t_{i}<t}y_{t_{i}}g(t-t_{i}))^{y_{t}}\exp({-\mu-\alpha\sum_{i:t_{i}<t}y_{t_{i}}g(t-t_{i})})}{y_{t}!}.$
The corresponding likelihood function is then,
$L(t)$ $\propto\prod_{t=1}^{T}\bigg{(}\mu+\alpha\sum_{i:t_{i}<t}y_{t_{i}}g(t-t_{i})\bigg{)}^{y_{t}}\exp(-\bigg{(}\mu+\alpha\sum_{i:t_{i}<t}y_{t_{i}}g(t-t_{i})\bigg{)}).$ (2)
1.2 Related work and contributions
@@
To our knowledge, we propose the first Bayesian, spatiotemporal Hawkes model applied to the ACLED data. We study countries in South Asia as an exemplar of how this model can be applied to these types of data worldwide. The models are fitted on a univariate level, so given sufficient computational resources the number of countries included in the study can be extended to include a much larger sample. Different types of conflict events are also considered individually, allowing us to compare the patterns of excitation for different conflict types. The proposed model enables a data-driven approach to monitoring and understanding the current risk of conflict events throughout time and space. Combined with input and expertise from key actors in the humanitarian and political sciences sectors, such as ACLED, this work can have important implications for the understanding and management of these events.
1.3 Outline
Section 2 introduces the data, the spatiotemporal model, and the inference methods used in this article. Section 3 follows by presenting the results of our analysis. We first present our Bayesian, spatiotemporal model, assess the model fit and interpret the outputs. Next, we present an alternative, less computationally intensive approach using maximum likelihood estimation. In Section 4, we demonstrate how our models can be used to assess conflict risk in practice and benchmark against the types of risk models used by organisations, such as ACLED, that rely on historical averages. Finally, Section 5 concludes with a discussion of the outcomes of this analysis, limitations of the model and avenues for future work.
@@ 2.1 Data
The ACLED project (Raleigh et al., 2010) provides a comprehensive geo-referenced database containing details of political violence and protests worldwide and has recorded in total more than one million individual events. In addition to the date and location of events, the database also records the number of fatalities, type of violence and the actors involved, which allows the user to further analyse and understand conflict events. Summaries and reports on the ACLED website are often segmented by conflict type, indicating that this type of segmentation is of interest to key stakeholders in the humanitarian sector. The types of conflicts recorded in the ACLED database and their respective definitions (ACLED, 2019) are,
@@
As mentioned in Section 1, for the purposes of illustrating our approach, we focus on South Asia as an exemplar region. We consider monthly counts of conflict events in South Asia over a 5 year period from 2010–2014. During the selected 5 year period, there were four countries with conflict data recorded in the ACLED database, namely Bangladesh, Sri Lanka, Nepal and Pakistan. Several other countries comprise the South Asian region, but did not have data available during the observation window for this study. Data from South and Southeast Asia start from 2010 in the ACLED dataset, and the 5 year period was chosen to reduce computational expense and because it avoids the period in 2016 when ACLED significantly increased capacity and updated their reporting methodology.
We are also interested in modelling spatial dynamics in these data to understand how conflict intensity is influenced by events in neighbouring regions. The inclusion of spatial information also enables the comparison of spatial dynamics between countries and conflict types. Despite having an additional parameter, we find that estimating a spatial kernel does not generally deteriorate the model fit based on the Bayesian Information Criterion (BIC). Further details are available in the supplementary materials.
@@
To determine the spatial aggregation we discretised according to administrative boundaries, which define the borders of a geographic area under a given political or social structure, rather than considering the individual location of each event. These boundaries were obtained from the Database of Global Administrative Areas (GADM) (Global Administrative Areas, 2012) and the Level 2 county level administrative boundaries were used in this analysis. The coordinates for each administrative boundary was then defined as the centroid of each region.
The data were aggregated to the monthly level, and broken down by each of the countries and conflict types present in our analyses. A daily temporal aggregation was considered initially; however, this was found to be unsuitable due to substantial zero-inflation present in the data at finer this discretisation, particularly when breaking down further into spatial regions. For completeness, this preliminary analysis is summarised in the supplementary materials.
@@ 2.2 Spatiotemporal discrete-time Hawkes process
The temporal DTHP defined in Section 1.1 can be extended to a spatiotemporal variation that incorporates a spatial triggering kernel. This means the level of excitation is then governed by both the time since an event elapsed and distance from the location where the event occurred. Incorporating spatial dependence into the existing model gives the following rate parameter $\lambda(t,x,y)$ for a particular time $t$ and location with coordinates $(x,y)$,
$\lambda(t,x,y)=\mu+\alpha\sum_{(t_{i},x_{i},y_{i}):t_{i}<t}y_{(t_{i},x_{i},y_{i})}g(t-t_{i})h(x-x_{i},y-y_{i})$ (3)
where $y_{(t,x,y)}$ denotes the number of events at time $t$ in location $(x,y)$ and $h(\cdot)$ is the spatial triggering kernel that determines how the distance between events affects the intensity of the process.
@@
Similar to the the temporal case, the number of events observed during time interval $t$ and location $(x,y)$, $Y(t,x,y)$ is assumed to be Poisson distributed such that $Y(t,x,y|\mathcal{H}_{t-1})\sim\text{Pois}(\lambda(t,x,y))$. If the set of centroids for all spatial regions is $\mathcal{S}\subset\mathbb{R}^{2}$, the probability that time period $t$ and location with coordinates $(x,y)$ has $y_{(t,x,y)}$ events is,
$L(t,x,y)$ $\propto\prod_{t=1}^{T}\prod_{(x,y)\in(\mathcal{X},\mathcal{Y})}\bigg{(}\mu+\alpha\sum_{(t_{i},x_{i},y_{i}):t_{i}<t}y_{(t_{i},x_{i},y_{i})}g(t-t_{i})h(x-x_{i},y-y_{i})\bigg{)}^{y_{(t,x,y)}}$
×$\times\exp(-\bigg{(}\mu+\alpha\sum_{(t_{i},x_{i},y_{i}):t_{i}<t}y_{(t_{i},x_{i},y_{i})}g(t-t_{i})h(x-x_{i},y-y_{i})\bigg{)}).$ (4)
The model parameters could be allowed to vary by spatial region. Moreover, we did consider a simple model for which $\mu$ and $\alpha$ varied based on a coarsely discretised population indicator. However, this simple approach had very little impact on the results, and further research is necessary to better understand the relationship between population size and the intensity of conflict events. Further details are available in the supplementary materials; however, in this article we will only consider the case where the model parameters are constant for all locations. Furthermore, allowing spatially-varying model parameters, in addition to issues with computational resources, could also result in poor model fit for regions with low event counts.
An important distinction between temporal and spatiotemporal DTHPs is in the interpretation of $\alpha$. The mass of excitation coming from another region $(x_{i},y_{i})$ is given by $\alpha h(x-x_{i},y-y_{i})$. Thus the global mass of excitation received from a previous time period $t_{i}$ is $\alpha$ multiplied by a sum over all locations of events that occurred at time $t_{i}$.
@@
As it is presented in (3), the temporal triggering kernel $g(\cdot)$ is defined on $\mathbb{Z}^{+}$. However, in many contexts its reasonable to assume that the maximum length of excitation is not infinite, and the impact of past events will eventually dissipate. Thus, the triggering kernel could be truncated to some finite value $t_{\max}$ with little impact when there is negligible mass at higher time intervals. Truncating the triggering kernel also reduces computational expense since the summations involved in (4) comprise less terms. In this study, we therefore specify the temporal triggering kernel in the form of a truncated, re-normalised geometric distribution such that
$g(s)=\frac{\beta(1-\beta)^{s-1}}{\sum_{j=1}^{t_{\max}}\beta(1-\beta)^{j-1}}.$
where $\beta\in[0,1]$ represents the success probability for the geometric distribution.
Furthermore, we represent the spatial triggering kernel $h(\cdot)$ by a radial basis function (RBF) kernel. Incorporating a spatial dimension into these models increases the computational cost, thus the RBF kernel was selected as it is simple and cheap to evaluate, and it is the most common spatial triggering kernel used for these models. For vectors of coordinates $z=(x,y)$ and $z_{i}=(x_{i},y_{i})$, the RBF spatial decay kernel $h$ is,
$h(z,z_{i})=\exp(-\frac{d(z,z_{i})}{2\sigma^{2}})$ (5)
@@
where $d$ is assumed to be the Euclidean distance function with bandwidth $\sigma$ such that,
$d(z,z_{i})=\sqrt{(x-x_{i})^{2}+(y-y_{i})^{2}}.$ (6)
@@
We note that distance in the latitude-longitude coordinate space can be poorly defined, as the latitude coordinate grid is equally spaced but the longitude has a strong dependence on the associated latitude value. This distorts the comparison of distances, thus if countries further away from the equator were being considered, then an alternative distance such as the Haversine distance should be used.
2.3 Inference

### v2

@@ Abstract
The monitoring of conflict risk in the humanitarian sector is largely based on simple historic averages. The overarching goal of this work is to assess the potential for using a more statistically rigorous approach to monitor the risk of political violence and conflict events in practice, and thereby improve our understanding of their temporal and spatial patterns, to inform preventative measures.
In particular, a Bayesian, spatiotemporal variant of the Hawkes process is fitted to data gathered by the Armed Conflict Location and Event Data (ACLED) project to obtain sub-national estimates of conflict risk in South Asia over time and space. Our model can effectively estimate the risk level of these events within a statistically sound framework, with a more precise understanding of uncertainty than was previously possible. The model also provides insights into differences in behaviours between countries and conflict types. We also show how our model can be used to monitor short and long term trends, and that it is more stable and robust to outliers compared to current practices that rely on historical averages.
@@ 1 Introduction
Improved modelling of conflict and political violence events has the potential to significantly advance several UN Sustainable Development Goals (SDGs). For SDG 16 (Peace, Justice and Strong Institutions), better modelling can help to identify drivers of instability, provide early warning signs, and highlight high-risk regions or populations that may need particular attention, supporting SDG 16’s targets on reducing violence, ensuring equal access to justice and building accountable institutions. It also contributes to SDG 1 (No Poverty) and SDG 2 (Zero Hunger), as conflict is a major cause of poverty and food insecurity. Furthermore, it has implications for SDG 3 (Good Health and Wellbeing) and SDG 4 (Quality Education), since conflict and violence can inhibit the effective functioning of health systems and educational institutions.
@@
Hawkes processes (Hawkes, 1971) are self-exciting stochastic processes whereby previous events in the process increase the probability of future events occurring. They have previously been used to study phenomena such as earthquakes (Ogata, 1988), crime and terrorism (White et al., 2013; Mohler et al., 2011), and social media interactions (Chen and Tan, 2018; Dutta et al., 2020). In this article, we extend the standard temporal model to include a spatially varying baseline rate of events, in addition to a spatial kernel that quantifies the spatial impact of events to the intensity of the process. A discrete-time variant of the spatiotemporal Hawkes process is fit to the ACLED data at a fine spatial scale to obtain sub-national estimates of conflict risk. This approach aids in characterising the temporal and spatial patterns that emerge from these types of conflicts, and provides an estimate of risk over space and time that can be used both for real-time monitoring and for making short-term predictions. We use data from the region of South Asia to showcase our approach; however, the methodology can easily be extended to other countries. For the purposes of this example, the four countries considered are Bangladesh, Sri Lanka, Nepal and Pakistan. In addition to modelling each of the countries separately, the various types of conflict events in each country are also modelled individually. By doing so, we can compare the spatial and temporal dynamics of events between both countries and conflict types.
@@
To our knowledge, we propose the first Bayesian, spatiotemporal Hawkes model applied to the ACLED data. We study countries in South Asia as an exemplar of how this model can be applied to these types of data worldwide. The models are fitted on a univariate level, so given sufficient computational resources the number of countries included in the study can be extended to include a much larger sample. Different types of conflict events are also considered individually, allowing us to compare the patterns of both temporal and spatial excitation for different conflict types. The proposed model enables a data-driven approach to monitoring and understanding the current risk of conflict events throughout time and space. Combined with input and expertise from key actors in the humanitarian and political sciences sectors, such as ACLED, this work can have important implications for the understanding and management of these events.
Section 2 introduces the data, the spatiotemporal models, and the inference methods used in this article. Section 3 follows by presenting the results of our analysis. We first present our Bayesian, spatiotemporal model, assess the model fit and interpret the outputs. In Section 4, we demonstrate how our models can be used to assess conflict risk in practice and benchmark against the types of risk models used by organisations, such as ACLED, that rely on historical averages. Finally, Section 5 concludes with a discussion of the outcomes of this analysis, limitations of the model and avenues for future work.
@@ 2.1 Data
The ACLED project (Raleigh et al., 2010) provides a comprehensive geo-referenced database containing details of political violence and protests worldwide and has recorded in total more than one million individual events. The earliest recorded events in the database are from 1997 for many countries in Africa, and the coverage has been progressively expanded over recent years to now having achieved global coverage with data being updated weekly. In addition to the date and location of events, the database also records the number of fatalities, type of violence and the actors involved, which allows the user to further analyse and understand conflict events. Summaries and reports on the ACLED website are often segmented by conflict type, indicating that this type of segmentation is of interest to key stakeholders in the humanitarian sector. The types of conflicts recorded in the ACLED database and their respective definitions (ACLED, 2019) are,
@@
As mentioned in Section 1, for the purposes of illustrating our approach, we focus on South Asia as an exemplar region. We consider weekly counts of conflict events in South Asia over a 5 year period from 2010–2014. During the selected 5 year period, there were four countries with conflict data recorded in the ACLED database, namely Bangladesh, Sri Lanka, Nepal and Pakistan. Data for South and Southeast Asia start from 2010 in the ACLED dataset, and the 5 year period was chosen to reduce computational expense and because it avoids the period in 2016 when ACLED significantly increased capacity and updated their reporting methodology. Several other countries comprise the South Asian region, but did not have data available during the observation window for this study.
Figure 1 shows the number of conflict events recorded each week for each conflict type and country over the observation window. There is a wide range of dynamics inherent in these data. We can see that Pakistan experienced substantially higher event counts over this period and Sri Lanka experienced the lowest number of conflict events. We also observe that some event types are more prevalent across all countries, such as protests and riots. Figure 2 displays the observed proportion of each conflict type for each country throughout the observation window. While protests and riots account for the majority of conflict events recorded, we see that Pakistan also has a high proportion of Explosions/remote violence compared to the other countries. We also see that Pakistan and Bangladesh experienced higher proportions of events classed as Battles, and that Sri Lanka and Bangladesh have higher proportions of events classed as Violence against civilians.
Figure 1: Weekly events in each country for each conflict type.
Figure 2: Proportion of total events of each conflict type by country.
@@
We are also interested in modelling spatial correlation in these data to understand how conflict intensity is influenced by events in neighbouring regions and to compare spatial dynamics between countries and conflict types. To determine the spatial aggregation we discretised according to administrative boundaries, which define the borders of a geographic area under a given political or social structure, rather than considering the individual location of each event. These boundaries were obtained from the Database of Global Administrative Areas (GADM) (Global Administrative Areas, 2012). For Nepal and Pakistan we used the Level 2 county/district level administrative boundaries. For Bangladesh and Sri Lanka, we used the coarser aggregation of Level 1 state/province level administrative boundaries since the Level 2 boundaries for these countries are very finely defined. The coordinates for each administrative boundary was then defined as the centroid of each region.
Figure 3 presents the total number of events of each conflict type to occur over the 5 year observation window at the administrative regions specified above for each country. The maps show that most activity generally occurs in or near the capital cities. Pakistan also has a high concentration of protests on the southeastern border near the biggest city and economic capital of Karachi. By modelling the spatial correlation in these data we can characterise how events from these regions influence neighbouring areas. There is also significant zero-inflation present in these weekly event counts. Figure 4 presents the proportion of weeks for each administrative region where the number of events for each conflict type was 0. In general, Sri Lanka and to a lesser extent Nepal have a higher proportion of zero event counts compared to Bangladesh and Pakistan.
Figure 3: Total number of events over 5 year observation window. Black crosses indicate capital cities (including Karachi in Pakistan as the biggest city and economic capital).
Figure 4: Proportion of weeks with zero events. Black crosses indicate capital cities (including Karachi in Pakistan as the biggest city and economic capital).
2.2 Discrete-time Hawkes processes
Hawkes processes (HPs) (Hawkes, 1971) were introduced to model phenomena that exhibit self-exciting behaviour whereby events occur in groups or clusters. HPs were originally introduced as a continuous-time process. However, since the data collected by the ACLED project is in the form of daily counts we adopt a discrete-time version of the HP (Browning et al., 2022, 2021; Linderman and Adams, 2015; Mohler, 2013), referred to in this article as the discrete-time Hawkes process (DTHP).
For a given DTHP, the number of events that occur at a particular time can be distributed according to any appropriate distribution defined on $\mathbb{Z}_{\geq 0}$ (Porter and White, 2012; Laub et al., 2025). In our analysis, we assume that events are distributed according to a Poisson distribution to approximate the continuous-time counterpart (Kirchner, 2016b). The DTHP is characterised by this distribution and its conditional intensity function which governs the rate at which events occur. The intensity function is comprised of two components: the baseline process and the self-exciting process. The baseline process describes the arrival rate of events that occur independently, and not as a result of other events. The self-exciting component on the other hand, describes the clustering behaviour of the process. The aspect of the self-exciting component that describes the impact that past events have on the current intensity of the process is referred to as the triggering kernel, and is a function of the time that has elapsed since an event occurred.
To state this formally, consider a linear DTHP $N$, where $N_{t}$ represents the number of events up to and including the $t$-th time interval and let the random variable $Y_{t}=N_{t}-N_{t-1}$ represent the number of event occurrences in $t$-th time interval. $N_{t}$ is then dependent on the set of events to occur up to but not including the $t$-th time interval, denoted by $\mathcal{H}_{t-1}=\{y_{s}:s<t\}$, where $y_{s}$ represents the observed number of events during the $s$-th time interval. We denote the corresponding conditional intensity function as $\lambda_{t}$. This quantity represents the number of events expected to occur in the counting process $N_{t}$ during the $t$-th time interval, conditionally on the past, and is given by,
$\lambda_{t}$ $=\mathbb{E}\{N_{t}-N_{t-1}|\mathcal{H}_{t-1}\}$
$=\mu+\alpha\sum_{s<t}y_{s}g(t-s)$ (1)
where $\mu$ represents the baseline mean of the process and the second term represents the self-exciting component of the DTHP. In particular, $\alpha$ is a magnitude parameter that describes the expected number of offspring events that a single event is expected to produce. If $\alpha>1$, then the dynamics of the process will be explosive and tend towards infinity, otherwise if $\alpha<1$ then the process will be stationary. The triggering kernel g(ts)g(t-s) is a probability mass function that describes the influence of these past events on the intensity of the process, given the time elapsed since the $s$-th time interval, where $s<t$.
The number of events observed during the $t$-th time interval, namely $y_{t}$, is distributed according to the random variable, $Y_{t}$, which has conditional mean $\mathbb{E}\{Y_{t}|\mathcal{H}_{t-1}\}=\lambda_{t}$ as defined in (1). In this analysis, $Y_{t}$ is assumed to be Poisson distributed such that $Y_{t}|\mathcal{H}_{t-1}\sim\text{Pois}(\lambda_{t})$. The corresponding likelihood function is then,
$L$ $=\prod_{t=1}^{T}\frac{\lambda_{t}^{y_{t}}}{y_{t}!}\exp(-\lambda_{t}).$ (2)
This model can be seen as a special case of a Poisson autoregressive model (Fokianos et al., 2009; Armillotta and Fokianos, 2024), and the endemic-epidemic model (Held et al., 2005; Bracher and Held, 2022; Lu and Meyer, 2023) which has been used widely in epidemiological research.
@@ 2.3 Spatiotemporal discrete-time Hawkes process
In this article, we consider several ways to extend the temporal DTHP defined in Section 2.2 to account for spatial patterns in the data. The first approach we consider is to include a spatially varying baseline rate $\mu_{x,y}$ for each region with coordinates $(x,y)$. Since we are aggregating to the level of administrative areas, this spatially varying parameter is also expected to indirectly capture underlying characteristics inherent in each region such as population size and other demographic factors. The second approach we consider is, in addition to the spatially varying baseline rate $\mu_{x,y}$, we also incorporate a spatial triggering kernel $h(\cdot)$ to capture any spatial clustering patterns that may be present. This means the level of excitation is then governed by both the time since an event elapsed and distance from the location where the event occurred. In both of these models we have also incorporated a term for zero-inflation to account for excess zeros in the data, as observed in Figure 4.
The first approach that includes a spatially varying baseline rate and no spatial self-excitation, referred to in this article as “nospatialSE”, results in the following rate parameter $\lambda_{t,x,y}$ for the $t$-th time interval and location with coordinates $(x,y)$,
$\lambda_{t,x,y}=\exp(\mu_{0}+\mu_{x,y})+\alpha\sum_{s<t}y_{s,x,y}g(t-s)$ (3)
where $\mu_{0}\in\mathbb{R}$ is a global baseline rate parameter, $\mu_{x,y}\in\mathbb{R}$ is a region-specific baseline rate, and their sum is exponentiated to ensure positivity. Furthermore, $y_{s,x,y}$ denotes the number of events to occur in the $s$-th time interval in location with coordinates $(x,y)$.
The second approach which incorporates a spatial triggering kernel in addition to the spatially varying baseline rate, referred to as “ST_SE”, gives the following rate parameter $\lambda_{t,x,y}$ for the $t$-th time interval and location with coordinates $(x,y)$,
$\lambda_{t,x,y}=\exp(\mu_{0}+\mu_{x,y})+\alpha\sum_{s<t}\sum_{(x_{l},y_{l})\in\mathcal{S}}y_{s,x_{l},y_{l}}g(t-s)h(x-x_{l},y-y_{l})$ (4)
where $\mathcal{S}$ is the set of centroids for all spatial regions, and $h(\cdot)$ is a spatial triggering kernel that determines how the distance between events affects the intensity of the process. Note also the subtle differences between (3) and (4), whereby the excitation term in the latter includes a summation over all other regions, and the former is only affected by events that have occurred in its own region.
@@
The likelihood function for both of these models is constructed in the same manner. We assume the number of events observed during $t$-th time interval in location with coordinates $(x,y)$, $Y_{t,x,y}$ follows a zero-inflated Poisson distributed such that $Y_{t,x,y}|\mathcal{H}_{t-1}\sim\text{Pois-ZeroInf}(\lambda_{t,x,y},\pi)$, where $\pi$ is the probability of a structural zero. The likelihood function is then,
$L$ $=\prod_{t=1}^{T}\prod_{(x,y)\in\mathcal{S}}\Big[\pi+(1-\pi)\exp(-\lambda_{t,x,y})\Big]^{\mathbb{I}\{y_{t,x,y}=0\}}\Big[(1-\pi)\dfrac{\lambda_{t,x,y}^{\,y_{t,x,y}}}{y_{t,x,y}!}\exp(-\lambda_{t,x,y})\Big]^{\mathbb{I}\{y_{t,x,y}\geq 1\}}$ (5)
An important distinction between temporal and spatiotemporal DTHPs as defined above is in the interpretation of $\alpha$. In the spatiotemporal setting under both models we propose, the incorporation of a zero-inflation factor means that the non-zero component of the likelihood function must be scaled by the probability of structural zeroes $\pi$, and as such $\alpha>1$ no longer necessarily corresponds to explosive dynamics. Additionally, when considering the model that includes a spatial triggering kernel as proposed in (4), the level of excitation from a single event also depends on the location of each event.
@@
As it is presented in (4), the temporal triggering kernel $g(\cdot)$ is defined on $\mathbb{Z}^{+}$. However, in many contexts it is reasonable to assume that the maximum length of excitation is not infinite, and the impact of past events will eventually dissipate. Thus, the triggering kernel could be truncated to some finite value $t_{\max}$ with little impact when there is negligible mass at higher time intervals. Truncating the triggering kernel also reduces computational expense since the summations involved in evaluating $\lambda_{t,x,y}$ comprise less terms.
We performed comparisons to assess three modelling assumptions, namely the maximum excitation time $t_{\max}$ (where we considered 12, 26 and 52 weeks), the form of $\lambda_{t,x,y}$, and the choice of triggering kernels $g(\cdot)$ and $h(\cdot)$. Across all event types and countries, we found that a longer $t_{\max}$ of 52 weeks provided the best fit to the data, and we therefore proceed with this value. Regarding the choice of model for $\lambda_{t,x,y}$, we considered three choices: no self-excitation (only a spatially varying baseline rate), a spatially varying baseline rate and only temporal self-excitation as in (3), and a spatially varying baseline rate with self-excitation both temporally and spatially as in (4). The first model with only a spatially varying baseline rate underperformed compared to the two alternatives. From these two models that include some form of self-excitation mechanism, neither emerged as a clear winner so both are considered in further analysis. Fitting both of these models for all country and conflict type scenarios also allows us to characterise more fully the spatial clustering dynamics by determining whether they are important in specific scenarios. In terms of triggering kernels, we compared a geometric temporal triggering kernel combined with a radial basis function (RBF) spatial triggering kernel against power law forms for both temporal and spatial kernels. There was no strong evidence supporting the inclusion of the more complex power law triggering kernels, as the additional parameters did not substantially improve model fit. We therefore retain the more parsimonious kernel specifications, which are described in detail below. Additional details on these comparisons are provided in the supplementary material.
In this study, we therefore specify the temporal triggering kernel in the form of a truncated, re-normalised geometric distribution such that,
$g(u)=\frac{\beta(1-\beta)^{u-1}}{\sum_{j=1}^{t_{\max}}\beta(1-\beta)^{j-1}}$ (6)
where $\beta\in[0,1]$ represents the success probability, or equivalently the inverse mean, for the geometric distribution. In the context of our study, a higher value of $\beta$ indicates that recent events have a stronger influence on the conditional intensity function $\lambda_{t,x,y}$. In other words, the impact of past events decays more rapidly over time, giving greater weight to events that occur closer to the current $t$-th time interval.
Furthermore, we represent the spatial triggering kernel by a radial basis function (RBF) kernel, which for vectors of coordinates $z=(x,y)$ and $z_{l}=(x_{l},y_{l})$ is given by,
$h(z,z_{l})=\exp(-\frac{d(z,z_{l})}{2\sigma^{2}})$ (7)
@@
where $\sigma$ is the bandwidth parameter and $d$ is assumed to be the Euclidean distance function such that,
$d(z,z_{l})=\sqrt{(x-x_{l})^{2}+(y-y_{l})^{2}}.$ (8)
In the context of our study, $\sigma$ determines how spatial proximity influences the conditional intensity function $\lambda_{t,x,y}$. A smaller value of $\sigma$ indicates that events occurring close to the region with coordinates $(x,y)$ have a stronger effect on $\lambda_{t,x,y}$, resulting in more localised spatial interactions. In contrast, a larger $\sigma$ implies a smoother spatial influence, whereby events occurring farther away can contribute more to the intensity of events at $(x,y)$.
@@
We also note that distance in the latitude-longitude coordinate space can be poorly defined, as the latitude coordinate grid is equally spaced but the longitude has a strong dependence on the associated latitude value. This distorts the comparison of distances, thus if countries further away from the equator were being considered, then an alternative distance such as the Haversine distance should be used.